OpenStudy (lukecrayonz):

Standard Deviation in statistics!!! please help:) The mean for stats test scores is 63 and the standard deviation is 6.0. For the bio test scores, the mean is 24 and the standard deviation is 3.8. The test scores of a student who took both tests are given below. A student gets a 76 on the stats test and a 29 on the bio test. a) Transform each test score into a z-score b) Determine on which test the student had a better score

OpenStudy (lukecrayonz):

@jim_thompson5910 :D

OpenStudy (lukecrayonz):

@satellite73

jimthompson5910 (jim_thompson5910):

hint: z = (x - mu)/sigma

jimthompson5910 (jim_thompson5910):

mu = mean sigma = standard deviation

jimthompson5910 (jim_thompson5910):

x is the raw score z is the z-score

OpenStudy (lukecrayonz):

so I got z=2 for the stats test

OpenStudy (lukecrayonz):

Not sure if that's off since it says round to the nearest hundredth

OpenStudy (lukecrayonz):

Nvm got 2.16 and 1.32

jimthompson5910 (jim_thompson5910):

stats test z = (x - mu)/sigma z = (76 - 63)/6 z = 2.1667

jimthompson5910 (jim_thompson5910):

So it would be 2.17 rounded to the nearest hundredth

jimthompson5910 (jim_thompson5910):

1.32 is correct for the bio test

OpenStudy (lukecrayonz):

Thank you, you're the only person who helps with my new stats questions:) Much appreciated

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (lukecrayonz):

@jim_thompson5910 one question

OpenStudy (lukecrayonz):

If the mean life of a tire is 32,000 miles, and a standard deviation of 2150, then 37000 would be unusual?

jimthompson5910 (jim_thompson5910):

what's the z score here

OpenStudy (lukecrayonz):

2.33

jimthompson5910 (jim_thompson5910):

Since P[Z>2.33] = 0.00990308, I'd say it's very unlikely to happen which means it's pretty unusual. Basically this score and any score above 2.33 is somewhat rare to occur. http://www.wolframalpha.com/input/?i=P%5BZ%3E2.33%5D